My grand daughter is about as many days as my son is in weeks, and my grand daughter is as many months as I am in years. My grand daughter, my son and I together are 120 years. Can you tell me my age in years

Answer by ankor@dixie-net.com(22740)


let m = his age
let s = son's age
let g = granddaughter's age
:
:
write an equation for each statement
:
My grand daughter is about as many days as my son is in weeks,
s = 7g
and my grand daughter is as many months as I am in years.
m = 12g
My grand daughter, my son and I together are 120 years.
g + s + m = 120
replace s with 7g; replace m with 12g
g + 7g + 12g = 120
20g = 120
g = 120/20
g = 6 yrs is granddaughter's age
:
Can you tell me my age in years
m = 12 * 6
m = 72 is your age
;
:
:
Check this; s = 7*6 = 42 is son's age
6 + 42 + 72 = 120